i need help.
write a system of equations.
an errand boy went to deposit some bills for his employer. some were one-dollar bills and the rest were five-dollar bills. the total value of the bills was $86 and the number of bills was 38. Find the number of each kind of bill.

Answer by htmentor(1343)

an errand boy went to deposit some bills for his employer. some were one-dollar bills and the rest were five-dollar bills. the total value of the bills was $86 and the number of bills was 38. Find the number of each kind of bill.
======================================
Let x = the number of five-dollar bills
Since the total number of bills is 38, the number of one-dollar bills will be 38-x
Then we can write the following equation for the total $ amount:
1*(38-x) + 5x = 86
Solve for x:
38 - x + 5x = 86
4x = 48
x = 12
Therefore, the number of 5-dollar bills is 12
And the number of 1-dollar bills is 38-12 = 26